# Exploring the Most Extreme Gamma-Ray Blazars Using Broadband Spectral Energy Distributions

### v1

@@ 3.1 Host galaxy emission
Since EHSP blazars have low non-thermal flux at optical/UV energies and their synchrotron peak occurs at higher frequencies, the host galaxy’s thermal emission is often prominent in the optical range of their SEDs. Therefore, before modelling the broadband SEDs, we fit different host galaxy models to the data in the optical range to determine the best-fitting model. For this purpose, we use four host galaxy templates from the SWIRE Template Library [^3] http://www.iasf-milano.inaf.it/~polletta/templates/swire_templates.html (Polletta et al., 2007). The host galaxies of most blazars are elliptical galaxies, characterised by old stellar populations and low star formation rates (Urry et al., 2000; Paliya et al., 2021). Lenticular galaxies, while distinct in structure, share these same characteristics and are indistinguishable from ellipticals in the optical band, specifically in the range where we fit the host galaxy templates ($7\times 10^{13}$ Hz to $10^{15}$ Hz). For this reason, the templates that we employ, shown in Figure 4, correspond to elliptical galaxies of 2 Gyr, 5 Gyr, and 13 Gyr, and to a lenticular galaxy. The best-fit host galaxy model is considered to be the one for which the lowest $\chi^{2}$ value is obtained in the optical range ($7\times 10^{13}$ Hz to $10^{15}$ Hz). We then assume black-body emission and, by applying Wien’s displacement law, estimate the approximate value of the temperature of the host galaxy. The host galaxy emission is generally not considered a significant source of seed photons for the external Compton process, and therefore, its contribution to the external Compton emission at high energies can be considered negligible. Thus, this contribution is not included in the broadband SED modelling. External Compton emission can also arise from the IR photons of the dusty torus. However, for EHSPs, this contribution is usually negligible, and, therefore, we exclude it from our modelling.
@@ 3.2 Synchrotron and inverse Compton emission
Once the host galaxy template that best fits the optical data is identified, the broadband SED modelling is conducted using a one-zone synchrotron/SSC model combined with the best-fit host galaxy model and all the data described in Section 2.2. In the absence of flaring states, the SEDs of HSPs and extreme blazars are well described by a simple one-zone SSC model (e.g., Sol & Zech, 2022), where the higher energy peak results from inverse Compton scattering of the electron population with photons produced by the synchrotron process. For these blazars, the external Compton component is typically negligible due to the relatively simple environments surrounding their jets, which lack a dusty torus or a BLR to supply the photons necessary for the external Compton mechanism (e.g., Ghisellini et al., 2011; Falomo et al., 2014). Therefore, a one-zone SSC model is generally sufficient to describe their SEDs (Celotti & Ghisellini, 2008).
@@ 3.3 Model assumptions and fitting
In the one-zone SSC model, both synchrotron and inverse Compton components are assumed to originate from the same region. For the source geometry, we adopt the simplest scenario where the emission is produced in a single spherical region or blob of radius $R$, located within the jet at a distance $R_{H}$ from the central SMBH. This region is filled with ultra-relativistic electrons moving with a bulk Lorentz factor $\Gamma$. In blazars, the observed emission is highly beamed and Doppler-boosted due to the blob’s relativistic motion and the jet’s small angle relative to our line of sight. The magnetic field $B$ is considered homogeneous within the blob. The electron population is modelled with a broken power-law distribution: a lower energy population with spectral slope $p_{1}$ and Lorentz factors between $\gamma_{\text{min}}$ and $\gamma_{\text{break}}$, and a higher energy population with spectral slope $p_{2}$ and Lorentz factors between $\gamma_{\text{break}}$ and $\gamma_{\text{max}}$. During the fitting, following the methodology from Nievas Rosillo et al. (2022), $\gamma_{\text{min}}$ is fixed at $10^{3}$, while $\gamma_{\text{max}}$, $\gamma_{\text{break}}$, $p_{1}$, and $p_{2}$ are treated as free parameters.
@@
The SED modelling is performed using the JetSeT v1.3.0 package (Tramacere, 2020; Tramacere et al., 2011, 2009), first running the lsb (least square bound scipy) minimiser followed by the minuit minimiser. To reduce the degrees of freedom in the model, some parameters are fixed to typical values for EHSP-like blazars, following Nievas Rosillo et al. (2022). The radius of the emitting region is set to $R=10^{16}$ cm, the distance of the blob from the central black hole is set to $R_{H}=2\times 10^{18}$ cm, and the bulk Lorentz factor is fixed at $\Gamma=20$. Adjusting these parameters impacts SED modelling: a larger emitting region lowers radiation energy density, affecting synchrotron and inverse Compton peaks; changing the SMBH distance alters jet dynamics and energy transfer; varying the Lorentz factor modifies observed flux through Doppler boosting. Despite these sensitivities, the parameters reflect standard EHSP models and general high-frequency blazar behaviour. With these assumptions, seven free parameters remain in the model: the magnetic field strength $B$, the electron density $N$, the spectral indices of the low- and high-energy particles $p_{1}$ and $p_{2}$, the maximum Lorentz factor $\gamma_{\text{max}}$, the break Lorentz factor $\gamma_{\text{break}}$, and the jet viewing angle $\theta$. The Doppler boosting factor $\delta$ is also treated as a free parameter, as it depends on the jet viewing angle $\theta$. Its dependence on the bulk Lorentz factor and the jet viewing angle is given by:
$\delta=\frac{1}{\Gamma(1-\beta\cos\theta)}\>,\quad\beta=v/c$ (1)
where $\beta$ is the speed of the jet in terms of the speed of light.
@@
We fit the data in the range from $5\times 10^{10}$ Hz to $10^{27}$ Hz, excluding the radio emission, as it is often believed to originate from a region much larger than the compact region responsible for the rest of the broadband SED. Additionally, at radio frequencies, synchrotron self-absorption becomes significant and may suppress low-frequency emission (e.g., Tavecchio et al., 1998). EBL absorption of $\gamma$ rays is taken into account, with attenuation becoming significant above $\sim$100 GeV for nearby sources but affecting progressively lower energies at higher redshifts, starting around $\sim$30 GeV for $z\sim 1$. The EBL, composed of the accumulated and redshifted light from all galaxies, stars, and interstellar dust in the universe, attenuates $\gamma$ rays through pair-production interactions with high-energy photons. This attenuation is incorporated using the optical depths from Domínguez et al. (2024b), based on the EBL model by Saldana-Lopez et al. (2021). An example of the broadband SED of one source from the sample fitted with this model is shown in Figure 5. All multi-wavelength SEDs of the 124 sample sources with their best-fit models are presented in Appendix A.
@@
Even after applying cuts to the initial sample to exclude potential variable sources in the X-ray band, when considering data from different instruments some sources still show flux dispersion. This variability in X-ray flux could have influenced the SED modelling. To ensure reliable results for the EHSP blazar properties derived here, we exclude sources with poor fitting results by selecting only those with $\chi^{2}/\text{dof}<1.5$. The $\chi^{2}$ was calculated using only the X-ray and $\gamma$-ray bands, considering all data points above $10^{16}$ Hz, since the radio and optical ranges show more dispersion. In total, 113 sources meet this requirement, resulting in the exclusion of 11 sources. These 11 sources are not considered when deriving the general properties of our blazar sample, as shown in Figures 67891011, and  12.
@@
Figure 5: Multi-wavelength emission of blazar 4FGL J0013.9--1854 modelled using a one-zone SSC model and a 13 Gyr old elliptical galaxy template for the host. The fluxes used to reconstruct the SEDs are shown as coloured points. The synchrotron and SSC emission from the best-fit model (dashed-grey line), the host galaxy emission (orange), the total intrinsic emission summing all components (solid-green line), and the observed emission after accounting for EBL absorption (solid-black line) are displayed.
@@ 4.1 Broadband SED modelling results
The best-fit parameters resulting from the SED modelling performed for each source in the sample are provided in Appendix B, Table LABEL:table:bestfit [^4] We make publicly available at Zenodo all the SED fluxes and the results from the SED fits (insert link) and at https://www.ucm.es/blazars/.. The SED modelling also allows us to calculate the synchrotron peak frequency, $\nu_{sync}^{peak}$, and the Compton dominance (CD), which are listed in Table LABEL:table:energy-budget for each source. CD is a parameter used in the study of blazars and other AGNs to describe the relative strength of inverse Compton emission compared to synchrotron emission in their spectral energy distributions. It plays a crucial role in understanding the physical environment of the jet, including the balance of magnetic and particle energy densities, the photon fields driving the inverse Compton process, and the overall energy dissipation mechanisms in the blazar (e.g., Paliya et al., 2021).
@@
For the jet viewing angle, we obtain values of $\theta<12^{\circ}$, with 85% of the sources having $\theta<8^{\circ}$. This corresponds to Doppler factors between $\sim 2$ and 40, typical for blazars due to their jets being closely aligned with our line of sight. For the maximum Lorentz factor $\gamma_{\text{max}}$, we find values ranging from $2.9\times 10^{5}$ to $4.6\times 10^{8}$. Determining $\gamma_{\text{max}}$ is challenging due to the lack of measurements between $10^{18}$ Hz and $\sim 10^{20}$ Hz, the gap between NuSTAR and Fermi-LAT energy ranges, which may result in $\gamma_{\text{max}}$ values larger than expected.
@@
Figure 6 shows the relationship between CD, magnetic field strength, and synchrotron peak frequency for the 113 sources with good fitting results. Sources with lower synchrotron peak frequencies generally have lower magnetic field values and higher CD, while the most extreme sources (with the highest synchrotron peak frequencies) display higher magnetic field values and lower CD. Figure 7 presents the CD distribution for the EHSP and HSP/ISP candidates in our sample, showing that EHSP candidates typically have lower CD values than less energetic sources, as demonstrated by Paliya et al. (2021). However, all sources in our sample have modest CD, with the highest being CD $=3.2$, corresponding to an HSP source (4FGL J1419.3$+$0444, $\nu_{sync}^{peak}=1.0\times 10^{16}$ Hz) based on our results. In contrast, the EHSP candidate with the highest CD value has CD $=1.5$ (4FGL J1719.3$+$1205).
@@
Figure 6: CD as a function of the magnetic field for the 113 sources in the sample with a good fitting result. The colour scale represents the synchrotron peak frequency.
@@
Figure 7: CD distribution of the sources in the sample with $\nu_{sync}^{peak}\geq 10^{17}$ Hz (EHSPs, in blue) and $\nu_{sync}^{peak}<10^{17}$ Hz (HSPs/ISPs, in orange).
@@
Figure 8: The CD, synchrotron peak frequency, and inverse Compton peak frequency obtained in this work are compared with those reported by Paliya et al. (2021) for the 28 overlapping sources. The synchrotron peak frequency $\nu_{sync}^{peak}$ and CD obtained in this work are also compared with the $\nu_{sync}^{peak}$ and CD values from Nievas Rosillo et al. (2022) for the 9 sources present in both samples. The dashed grey line indicates the one-to-one correspondence.
@@
Out of the 113 sources with good fits, 28 overlap with those in Paliya et al. (2021), which catalogues 1077 $\gamma$-ray blazars or BCUs detected by Fermi-LAT with available optical spectra or measurements of host galaxy bulge magnitude, central SMBH mass, or disk luminosity. Figure 8 compares our results for CD and the synchrotron ($\nu_{sync}^{peak}$) and inverse Compton ($\nu_{IC}^{peak}$) peak frequencies for these 28 sources with those from Paliya et al. (2021). Their model does not account for host galaxy emission, potentially increasing the modelled non-thermal synchrotron component. This could partly explain why our $\nu_{sync}^{peak}$ values are systematically higher by a factor around 6, although a small contribution may arise from the inherent uncertainty introduced by the manual fitting procedure and differences between the criteria adopted in each work, the data selection or the modelling codes.
@@
Furthermore, a correlation between accretion luminosity ($L_{\text{disk}}$) and CD in blazars was found by Paliya et al. (2021, shown in Figure 11 of that work), suggesting that blazars with high CD tend to be more luminous. In particular, emission-line blazars generally have CD $>1$ and $L_{\text{disk}}/L_{\text{Edd}}>0.01$, while absorption-line blazars typically have CD $<1$ and $L_{\text{disk}}/L_{\text{Edd}}<0.01$. Based on these findings and given that most of the EHSP candidates in our sample show CD values of $\text{CD}\lesssim 1$, we expect these blazars to generally have $L_{\text{disk}}$ values in Eddington units of $L_{\text{disk}}/L_{\text{Edd}}\leq 0.01$. Therefore, they can be classified as low-Compton-dominated (LCD) objects, according to the classification proposed by Paliya et al. (2021). In contrast, more Compton-dominated blazars, primarily FSRQs, are typically classified as high-Compton-dominated (HCD) objects. Hence, we conclude that our blazar sample mainly consists of LCD objects, with no emission lines and low accretion activity.
@@
Out of the 113 sample sources with good fitting results, 28 are classified as BCUs in the 4FGL catalogue. According to our modelling results, all but two of these BCUs have CD $<1$, showing similar emission properties to those of BL Lacs. The remaining two BCUs, for which we obtain CD $\geq 1$, are 4FGL J0611.1+4325 (with CD $=1.4$) and 4FGL J1719.3+1205 (with CD $=1.5$).
An analysis of potential correlations between black hole mass and key jet properties, including jet viewing angle, magnetic field strength, energy density ratios, jet power components, synchrotron peak frequency, and CD, for the sources shared with Paliya et al. (2021) reveals no significant trends. This suggests that jet emission in EHSPs is primarily governed by local jet conditions rather than the mass of the central black hole. Moreover, EHSPs consistently show extreme synchrotron peaks and low CD across the full range of black hole masses.

### v2

@@ 3.1 Host galaxy emission
Since EHSP blazars have low non-thermal flux at optical/UV energies and their synchrotron peak occurs at higher frequencies, the host galaxy’s thermal emission is often prominent in the optical range of their SEDs. Therefore, before modelling the broadband SEDs, we fitted different host galaxy models to the data in the optical range to determine the best-fitting model. For this purpose, we used four host galaxy templates from the SWIRE Template Library [^3] http://www.iasf-milano.inaf.it/~polletta/templates/swire_templates.html (Polletta et al., 2007). The host galaxies of most blazars are elliptical galaxies, characterised by old stellar populations and low star formation rates (Urry et al., 2000; Paliya et al., 2021). Lenticular galaxies, while distinct in structure, share these same characteristics and are indistinguishable from ellipticals in the optical band, specifically in the range where we fit the host galaxy templates ($7\times 10^{13}$ Hz to $10^{15}$ Hz). For this reason, the templates that we employed, shown in Figure 4, correspond to elliptical galaxies of 2 Gyr, 5 Gyr, and 13 Gyr, and to a lenticular galaxy. The best-fit host galaxy model was considered to be the one for which the lowest $\chi^{2}$ value is obtained in the optical range ($7\times 10^{13}$ Hz to $10^{15}$ Hz). We then assumed black-body emission and, by applying Wien’s displacement law, estimated the approximate value of the temperature of the host galaxy. The host galaxy emission is generally not considered a significant source of seed photons for the external Compton process, and therefore, its contribution to the external Compton emission at high energies can be considered negligible. Thus, this contribution was not included in the broadband SED modelling. External Compton emission can also arise from the IR photons of the dusty torus. However, for EHSPs, this contribution is usually negligible, and, therefore, we excluded it from our modelling.
@@ 3.2 Synchrotron and inverse Compton emission
Once the host galaxy template that best fits the optical data was identified, the broadband SED modelling was conducted using a one-zone synchrotron/SSC model combined with the best-fit host galaxy model and all the data described in Section 2.2. In the absence of flaring states, the SEDs of HSPs and extreme blazars are well described by a simple one-zone SSC model (e.g. Sol & Zech, 2022), where the higher energy peak results from inverse Compton scattering of the electron population with photons produced by the synchrotron process. For these blazars, the external Compton component is typically negligible due to the relatively simple environments surrounding their jets, which lack a dusty torus or a BLR to supply the photons necessary for the external Compton mechanism (e.g. Ghisellini et al., 2011; Falomo et al., 2014). Therefore, a one-zone SSC model is generally sufficient to describe their SEDs (Celotti & Ghisellini, 2008).
@@ 3.3 Model assumptions and fitting
In the one-zone SSC model, both synchrotron and inverse Compton components are assumed to originate from the same region. For the source geometry, we adopted the simplest scenario where the emission is produced in a single spherical region or blob of radius $R$, located within the jet at a distance $R_{H}$ from the central SMBH. This region is filled with ultra-relativistic electrons moving with a bulk Lorentz factor $\Gamma$. In blazars, the observed emission is highly beamed and Doppler-boosted due to the blob’s relativistic motion and the jet’s small angle relative to our line of sight. The magnetic field, $B$, is considered homogeneous within the blob.
The electron population was modelled with a smooth broken power law distribution: a lower energy population with spectral slope $p_{1}$ and Lorentz factors between $\gamma_{\text{min}}$ and $\gamma_{\text{break}}$, and a higher energy population with spectral slope $p_{2}$ and Lorentz factors between $\gamma_{\text{break}}$ and $\gamma_{\text{max}}$. The smooth broken power law provides an optimal balance between model flexibility and physical interpretability, making it a widely used framework for characterising HSPs and EHSPs (e.g. Aleksić et al., 2015; Sinha et al., 2017). To test the robustness of this choice, we applied two alternative electron energy distributions, i.e. a power law with an exponential cut-off and a log-parabola, to a subsample of the sources. Both alternatives yield results consistent with those obtained using the smooth broken power law, indicating that the choice of this electron energy distribution does not introduce significant bias in the modelling. During the fitting, following the methodology from Nievas Rosillo et al. (2022), $\gamma_{\text{min}}$ was fixed at $10^{3}$, while $\gamma_{\text{max}}$, $\gamma_{\text{break}}$, $p_{1}$, and $p_{2}$ were treated as free parameters. This approach is commonly adopted in modelling studies when low-energy data are ambiguous (e.g. Abe et al., 2023; Ouyang et al., 2025). To assess the impact of this choice, we repeated the SED modelling on a subsample of sources with $\gamma_{\text{min}}$ left as a free parameter. A comparison of the results obtained with fixed and free $\gamma_{\text{min}}$ reveals no systematic differences, confirming that fixing $\gamma_{\text{min}}$ to $10^{3}$ does not introduce significant bias into the model.
@@
The SED modelling was performed using the JetSeT v1.3.0 package (Tramacere, 2020; Tramacere et al., 2011, 2009), first running the lsb (least square bound) minimiser followed by the minuit minimiser. To reduce the degrees of freedom in the model, some parameters were fixed to typical values for EHSP-like blazars, following Nievas Rosillo et al. (2022). The radius of the emitting region was set to $R=10^{16}$ cm, the distance of the blob from the central black hole was set to $R_{H}=2\times 10^{18}$ cm, and the bulk Lorentz factor was fixed at $\Gamma=20$. Adjusting these parameters impacts SED modelling: a larger emitting region lowers radiation energy density, affecting synchrotron and inverse Compton peaks; changing the SMBH distance alters jet dynamics and energy transfer; varying the Lorentz factor modifies observed flux through Doppler boosting. Despite these sensitivities, the parameters reflect standard EHSP models and general high-frequency blazar behaviour. With these assumptions, seven free parameters remained in the model: the magnetic field strength, $B$; the electron density, $N$; the spectral indices of the low- and high-energy particles, $p_{1}$ and $p_{2}$; the maximum Lorentz factor , $\gamma_{\text{max}}$; the break Lorentz factor , $\gamma_{\text{break}}$; and the jet viewing angle , $\theta$. The Doppler boosting factor, $\delta$, was also treated as a free parameter, as it depends on the jet viewing angle, $\theta$. Its dependence on the bulk Lorentz factor and the jet viewing angle is given by
$\delta=\frac{1}{\Gamma(1-\beta\cos\theta)}\>,\quad\beta=v/c\>,$ (1)
where $\beta$ is the speed of the jet in terms of the speed of light. The parameters used in the SED modelling are listed in Table 1, along with their corresponding fit ranges and whether they were treated as free or frozen during the fitting process.
Table 1: Parameters used in the SED modelling.
Parameter Free/Frozen Fit Range/Value
R frozen $10^{16}$ cm
[…]
$\Gamma$ frozen 20
[…]
$\gamma_{\text{min}}$ frozen $10^{3}$
[…]
 [^4] The table lists whether each parameter was frozen or free during the fit, and the corresponding range over which fitting was performed.
@@
We fitted the data in the range from $5\times 10^{10}$ Hz to $10^{27}$ Hz, excluding the radio emission, as it is often believed to originate from a region much larger than the compact region responsible for the rest of the broadband SED. Additionally, at radio frequencies, synchrotron self-absorption becomes significant and may suppress low-frequency emission (e.g. Tavecchio et al., 1998). The absorption of $\gamma$ rays by the EBL was taken into account, with attenuation becoming significant above $\sim$100 GeV for nearby sources but affecting progressively lower energies at higher redshifts, starting around $\sim$30 GeV for $z\sim 1$. The EBL, composed of the accumulated and redshifted light from all galaxies, stars, and interstellar dust in the universe, attenuates $\gamma$ rays through pair-production interactions with high-energy photons. This attenuation was incorporated using the optical depths from Dom´ınguez et al. (2024b), based on the EBL model by Saldana-Lopez et al. (2021). An example of the broadband SED of one source from the sample fitted with this model is shown in Figure 5. All multi-wavelength SEDs of the 124 sample sources with their best-fit models are presented in Appendix A.
@@
Even after applying cuts to the initial sample to exclude potential variable sources in the X-ray band, when considering data from different instruments some sources still show flux dispersion. This variability in X-ray flux could have influenced the SED modelling. To ensure reliable results for the EHSP blazar properties derived here, we excluded sources with poor fitting results by selecting only those with $\chi^{2}/\text{dof}<1.5$. The $\chi^{2}$ was calculated using only the X-ray and $\gamma$-ray bands, considering all data points above $10^{16}$ Hz, since the radio and optical ranges show more dispersion. In total, 113 sources meet this requirement, resulting in the exclusion of 11 sources. These 11 sources were not considered when deriving the general properties of our blazar sample, as shown in Figures 67891011, and  12. In addition, to assess the potential bias introduced by limited X-ray coverage, we divided the full sample into two subsamples based on whether the obtained synchrotron peak frequency lay above the highest-energy available X-ray data. The comparison of their key properties revealed no systematic differences (aside from the expected shift in $\nu_{sync}^{peak}$), indicating that the absence of X-ray constraints on the peak position does not introduce a significant bias in our results.
@@
Figure 5: Multi-wavelength emission of blazar 4FGL J0013.9--1854 modelled using a one-zone SSC model and a 13 Gyr old elliptical galaxy template for the host. The fluxes used to reconstruct the SEDs are shown as coloured points. The synchrotron and SSC emission from the best-fit model (dashed grey line), the host galaxy emission (orange), the total intrinsic emission summing all components (solid green line), and the observed emission after accounting for EBL absorption (solid black line) are displayed.
@@ 4.1 Broadband SED modelling results
The best-fit parameters resulting from the SED modelling performed for each source in the sample are provided in Appendix B, Table LABEL:table:bestfit [^5] We make publicly available at Zenodo all the SED fluxes and the results from the SED fits (doi: 10.5281/zenodo.15778399) and at https://www.ucm.es/blazars/ehsp. See Appendix C for a description of the structure of the FITS file.. The SED modelling also allowed us to calculate the synchrotron peak frequency, $\nu_{sync}^{peak}$, and the Compton dominance (CD), which are listed in Table LABEL:table:energy-budget for each source. CD is a parameter used in the study of blazars and other AGNs to describe the relative strength of inverse Compton emission compared to synchrotron emission in their spectral energy distributions. It plays a crucial role in understanding the physical environment of the jet, including the balance of magnetic and particle energy densities, the photon fields driving the inverse Compton process, and the overall energy dissipation mechanisms in the blazar (e.g. Paliya et al., 2021).
@@
For the jet viewing angle, we obtain values of $\theta<12^{\circ}$, with 85% of the sources having $\theta<8^{\circ}$. This corresponds to Doppler factors between 2$\sim$ 2∼ 2 and 40, typical for blazars due to their jets being closely aligned with our line of sight. For the maximum Lorentz factor $\gamma_{\text{max}}$, we find values ranging from $2.9\times 10^{5}$ to $4.6\times 10^{8}$. Determining $\gamma_{\text{max}}$ is challenging due to the lack of measurements between $10^{18}$ Hz and $\sim 10^{20}$ Hz, the gap between NuSTAR and Fermi-LAT energy ranges, which may result in $\gamma_{\text{max}}$ values larger than expected.
@@
Figure 6 shows the relationship between CD, magnetic field strength, and synchrotron peak frequency for the 113 sources with good fitting results. Sources with lower synchrotron peak frequencies generally have lower magnetic field values and higher CD, while the most extreme sources (with the highest synchrotron peak frequencies) display higher magnetic field values and lower CD. Figure 7 presents the CD distribution for the EHSP and HSP/ISP candidates in our sample, showing that EHSP candidates typically have lower CD values than less energetic sources, as demonstrated by Paliya et al. (2021). However, all sources in our sample have modest CD, with the highest being CD =3.2=3.2= 3.2, corresponding to an HSP source (4FGL J1419.3$+$0444, $\nu_{sync}^{peak}=1.0\times 10^{16}$ Hz) based on our results. In contrast, the EHSP candidate with the highest CD value has CD =1.5=1.5= 1.5 (4FGL J1719.3$+$1205).
@@
Figure 6: Compton dominance as a function of the magnetic field for the 113 sources in the sample with a good fitting result. The colour scale represents the synchrotron peak frequency.
@@
Figure 7: Compton dominance distribution of the sources in the sample with $\nu_{sync}^{peak}\geq 10^{17}$ Hz (EHSPs, in blue) and $\nu_{sync}^{peak}<10^{17}$ Hz (HSPs/ISPs, in orange).
@@
Figure 8: Comparison of the CD, synchrotron peak frequency, and inverse Compton peak frequency obtained in this work with those reported by Paliya et al. (2021) for the 28 overlapping sources. The synchrotron peak frequency $\nu_{sync}^{peak}$ and CD obtained in this work are also compared with the $\nu_{sync}^{peak}$ and CD values from Nievas Rosillo et al. (2022) for the nine sources present in both samples. The dashed grey line indicates the one-to-one correspondence.
@@
Out of the 113 sources with good fits, 28 overlap with those in Paliya et al. (2021), which catalogues 1077 $\gamma$-ray blazars or BCUs detected by Fermi-LAT with available optical spectra or measurements of host galaxy bulge magnitude, central SMBH mass, or disc luminosity. Figure 8 compares our results for CD and the synchrotron ($\nu_{sync}^{peak}$) and inverse Compton ($\nu_{IC}^{peak}$) peak frequencies for these 28 sources with those from Paliya et al. (2021). Their model does not account for host galaxy emission, potentially increasing the modelled non-thermal synchrotron component. This could partly explain why our $\nu_{sync}^{peak}$ values are systematically higher by a factor around 6, although a small contribution may arise from the inherent uncertainty introduced by the manual fitting procedure and differences between the criteria adopted in each work, the data selection or the modelling codes.
@@
[…]
@@
Out of the 113 sample sources with good fitting results, 28 are classified as BCUs in the 4FGL catalogue. According to our modelling results, all but two of these BCUs have CD <1<1< 1, showing similar emission properties to those of BL Lacs. The remaining two BCUs, for which we obtain CD 1$\geq$ 1≥ 1, are 4FGL J0611.1+4325 (with CD =1.4=1.4= 1.4) and 4FGL J1719.3+1205 (with CD =1.5=1.5= 1.5).
An analysis of potential correlations between black hole mass and key jet properties, including jet viewing angle, magnetic field strength, energy density ratios, jet power components, synchrotron peak frequency, and CD, for the sources shared with Paliya et al. (2021) revealed no significant trends. This suggests that jet emission in EHSPs is primarily governed by local jet conditions rather than the mass of the central black hole. Moreover, EHSPs consistently show extreme synchrotron peaks and low CD across the full range of black hole masses.